Consider a series of $113 end-of-period CFs spanning 2050-2056 and a $50 CF at the beginning of 2042. The interest rate is 2.5%. What is the equivalent value of these CFs at the end of 2035?
Answer #1

To find the value of cash flows at the end of 2035, we will have to discount the cash flows generated in future

Hence, value of cash flow to be received at the beginning of 2042 or end of 2041 is 50*PVF(2.5%,6 Years)

= 50*0.862

=$43.1

Value of cash flows to be received from 2050-56 at the beginning of 2050 = 113*PVAF(2.5%, 7 Years)

= 113*6.349

=$717.437

Value at the end of 2035 = 717.437*PVF(2.5%, 14 Years)

= 717.437*0.708

=$507.95

Hence, total value of cash flows today= $551.05